Fermi National Laboratory


Accelerator Update

Monday June 7
The day shift began with Operations monitoring store 3560 and the Recycler stash, with the Antiproton Source (Pbar) stacking, and with MiniBooNE taking beam.

At 10:11 AM, Operations reported that there had been two beam spikes around 9 AM. Those spikes seem to have caused equipment to trip off at the D-zero (D0) experiment and perhaps the Collider Detector at Fermilab (CDF) as well.

At 10:58 AM, a Pbar Accumulator vacuum pump tripped off. Operators reset the pump.

Operations terminated store 3560 at 12:08 PM.

Operations put the Tevatron (TeV) through a dry squeeze at 12:25 PM and then at 12:50 PM put the TeV into shot setup.

Operations established store 3562 at 3:19 PM with an initial luminosity of 76.8446E30. This set another RECORD initial luminosity.

Main Injector (MI) and Recycler experts completed conducting Mixed Mode pilot transfers at 3:32 PM.

Operations tried to put Pbar back into its stacking mode at 3:44 PM, but problems developed. Operators investigated and discovered a bad restore file had been loaded. Stacking resumed at 3:56 PM.

Operation began sending beam to the Meson target train at 4:35 PM.

Recycler completed stashing antiprotons from Pbar at 7:44 PM. Pbar resumed stacking.

Operations established beam to Meson MCenter experiment at 9:11 PM.

Operators adjusted the TeV collimators and made tune changes to fix a colliding beam problem for CDF.

Tuesday June 8
The midnight shift began with Operations monitoring store 3562 and the stash, with Pbar stacking, and with MiniBooNE and MCenter taking beam.

At 12:57 AM, a Pbar Accumulator vacuum pump tripped off again. Operators left the pump off.

Operations reported at 5:32 AM that the TeV colliding beam suffered from another mysterious beam spike. Investigations showed that there was no earthquake activity to coincide with the spike. (Yes, the Tevatron acts as a very sensitive seismic monitor for both North and South America.)

CDF called about TeV colliding beam problems at 6:19 AM. Operators adjusted the dampers.

Operations reported a power glitch at 9:19 AM. Many devices all around the complex tripped off. The high voltage expert said there was an offsite trip on an incoming ComEd line.

MI experts began a one-hour MECAR study at 10:17 AM.

Pbar resumed stacking at 10:34 AM.

MI experts began a multi-batching study at 11:37 AM.

Operations reported at 11:53 AM that operators had been trying to resolve some colliding beam problems. They would make headway in one area only to cause more problems in other areas. An expert noticed that a CDF air conditioning system was off and that was the root cause.

MI experts began a slipstacking study at 1:58 PM.

Pbar experts began Debuncher and Accumulator studies at 3:05 PM.

Both MI and Pbar studies finished around 5 PM. Stacking resumed at 5:19 PM.

Wednesday June 9
The midnight shift began with Operations monitoring store 3562 and the stash, with Pbar stacking, and with MiniBooNE and MCenter taking beam.

At 5:08 AM, Operations reported that Pbar wasn't getting any beam. A MI kicker wasn't charging. Operators informed an expert.

Operations terminated store 3562 at 5:32 AM.

Pbar resumed stacking at 6:17 AM.

Plans for Today
The plan for this morning is the Mixed Mode transfer of antiprotons from the Pbar stack and the Recycler stash to the TeV for the next store.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 6/9/2004   email Fermilab